# Finding Safety Neurons in Large Language Models

### v1

@@ 6 Related work
Preference Learning. With the success of ChatGPT (OpenAI, 2023), aligning LLMs with human values and preferences—known as preference learning—has emerged as a key research focus. The Reinforcement Learning from Human Feedback (RLHF) paradigm, utilized in ChatGPT, becomes the dominant approach in this field (Bai et al., 2022a). However, due to the instability nature of reinforcement learning and the high resource consumption of RLHF training, various alternatives have been proposed, such as DPO (Rafailov et al., 2024), KTO (Ethayarajh et al., 2024), and SPPO (Wu et al., 2024). In this work, we focus on DPO-based alignment algorithms due to their simplicity and effectiveness, which have led to widespread adoption. Recent efforts have extended preference learning to areas such as reasoning Wang et al. (2023); Lambert et al. (2024) and information extraction Qi et al. (2024), showing promising results. Although our primary focus is on safety alignment, our method can be applied to other types of alignment without modification.
Mechanistic Interpretability for Transformer. Identifying interpretable neurons has long been a goal of mechanistic interpretability research in Transformers (Geva et al., 2021; Elhage et al., 2022; Gurnee et al., 2023, 2024). Geva et al. (2021) proposed viewing the feed-forward networks in Transformers as key-value memories, providing a new direction for interpretation. Dai et al. (2022) identified knowledge neurons through knowledge attribution, showing that their activations are positively correlated with the expression of corresponding facts. Wang et al. (2022b) discovered skill neurons within pre-trained Transformers, which are highly predictive of task labels, by computing their predictive scores for task labels. However, these methods are limited to tasks with few token labels and thus cannot be directly applied to safety alignment. Recent work (Lee et al., 2024) provides a mechanistic interpretation for DPO, while their experiments are limited to GPT-2 and detoxifying. In this work, we study general safety alignment on recent LLMs.
@@ 7 Conclusion
Limitations
Ethical Consideration
This work is devoted to exploring the underlying mechanisms of safety alignment—a critical technique to ensure the safety of LLMs. We aim to provide insights that will help the community develop safer applications using LLMs. We discuss the intended usage, potential misuse, and measures for risk control.
Intellectual property. All the datasets we used are open-sourced, and we strictly adhere to their licenses. We believe all the datasets are well-desensitized. For the investigated LLMs, we query GPT-4 through paid APIs. For Llama2111https://ai.meta.com/llama/license/, Mistral [^2] https://github.com/openstack/mistral/blob/master/LICENSE, and Gemma [^3] https://github.com/google-deepmind/gemma/blob/main/LICENSE we strictly adhere to their license. We obtain the Llama2’s checkpoint by applying to Facebook [^4] https://github.com/facebookresearch/llama.
Intended Usage. We designed a demonstrating technology to help prevent LLMs from generating harmful content, as demonstrated in Section 5. Furthermore, we encourage researchers to use our findings to monitor and correct misbehavior in LLMs. It is our hope that this paper will inspire the development of more robust technologies that better align LLMs with human values.
Potential Misuse. It is important to note the possibility of developing adversarial techniques that compromise safety by preserving safety neurons, potentially giving rise to more covertly malicious LLMs. Recognizing and mitigating this threat is crucial to maintaining the integrity and safety of LLM applications.
Risk Control. To mitigate potential risks, we will release our code and the data used in this paper. We believe that transparency will help reduce the risks associated with our work and facilitate the responsible use and further development of the technologies discussed.
@@ A.3 Evaluation Benchmarks
RedTeam -- $38,961$
HarmBench -- $400$
JailbreakLLMs -- $390$
LIMA -- $1,030$
@@
Table 4: Data statistics of the used datasets.
@@ B.1 Safety Alignment
Table 5: Hyperparameter used for SFT.
DPO Training Details
B.2 Safety Evaluation
B.3 Perplexity Evaluation
We evaluate the perplexity on the full test set of Wikitext-2 with a max length of 4096. We run all the above experiments on NVIDIA A100-SXM4-80GB GPU, and it takes about 1,000 GPU hours.
Model $d_{\mathrm{vocab}}$ $d_{\mathrm{model}}$ $d_{\mathrm{mlp}}$ $n_{\mathrm{layers}}$ $n_{\mathrm{heads}}$Neurons Activation
@@ B.3 Perplexity Evaluation
Table 6: Hyperparameter of LLMs studied.
@@ 5. Put a fake mustache or beard on
B.4 Finding Safety Neurons
We build our code on TransformerLens (Nanda and Bloom, 2022) to cache neuron activations and perform dynamic activation patching. For each prompt dataset, we use 200 randomly sampled prompts (no overlap with evaluation data). Again, we use greedy search for generation and set the max new tokens to 256, resulting in around 40,000 activations for each neuron.
B.5 Harmful Content Prediction
Appendix C More Properties of Safety Neurons
C.1 Layer Distribution
The layer distribution of these neurons is shown in Figure 6. Llama2-7b and Mistral-7b have similar patterns: safety neurons are distributed across many layers, predominantly appearing in the deep layers, with a gradual shift towards the middle layers as change scores decrease. Conversely, Gemma-7b presents a starkly different distribution, with safety neurons primarily found in the initial and final layers. Notably, the most significant neurons in Gemma-7b are located in shallower layers, progressively transitioning to deeper layers with a more uniform distribution as change scores decrease. This phenomenon is likely due to significant architectural differences between Gemma-7b and the other two models (Table 6).
Figure 6: The layer distribution of (20,000) safety neurons, grouped by every 5,000 neurons. The layer depth is the normalized layer number.
C.2 Change Score Distribution
We visualize the change scores distribution of safety neurons in Figure 7. We first notice that only a small fraction of neurons changed much after safety alignment (for Llama2-7b only 876 out of 341248 neurons with a change score larger than $0.1$). More interestingly, these three different models have similar patterns and thresholds at around $0.035$ for safety neurons. Furthermore, we find that models performing better in safety alignment exhibit longer tails [^5] The skewness of Llama2-7b, Mistral-7b-v0.1 and Gemma-7b are $6.99$, $7.20$ and $19.89$ respectively., indicating that improved model performance may result from more neurons experiencing significant activation changes. We leave the further investigation of this phenomenon for future work.
Figure 7: The distribution of change scores of (20,000) safety neurons (truncated for better visualization).
C.3 Change Scores are Appropriate Indicator of Safety Neurons
To further validate the change score as an appropriate indicator of neurons’ causal effect on generation, we conducted experiments using the same number of neurons but varying the change score ranges. Specifically, we used consecutive sets of 20,000 neurons, starting from different ranks. As depicted in Figure 8, we observed that as the change scores of the neurons decreased, the effectiveness of dynamic activation patching rapidly diminished. This result indicates that only neurons with high change scores have a significant causal effect on the model’s output.
Figure 8: Cost score of Base and SFT patched with different consecutive 20000 neurons. The horizontal axis represents the rank of the highest-ranked neuron among the 20000 neurons.
C.4 Specificity on Different Datasets
We simply use safety neurons found on HH-Harmless in previous experiments. Now we take a closer look at the prompt dataset selection. We use datasets from 3 different preference learning tasks: (1) Safety, including Beavertails (Ji et al., 2024), HH-Harmless (Bai et al., 2022a), and JailBreakLLMs (Shen et al., 2023); (2) Helpfulness, including HH-Harmless (Bai et al., 2022a) and LIMA (Zhou et al., 2024); (3) Reasoning, including the Reasoning subset from RewardBench (Lambert et al., 2024). We repeat the experiments from §3.1 using safety neurons found on these prompts, as shown in Figure 14. The results indicate that safety neuron activations are specific to certain inputs; safety neurons found on similar types of prompts exhibit similar causal effects and are most effective on safety-related prompts.
Appendix D Other Design Choices for Neuron-Finding
D.1 Which Model Should be Compared?
We explore the impact of comparing different models and different generations. We replicate the experiments from §3.1 with different design choices, and the results are depicted in Figure 15. More detailed results on various red-teaming benchmarks are presented in Table 9. Figure 15 demonstrates that the choice of models and generations for comparison does not fundamentally affect the method’s effectiveness. Table 9 further indicates that the optimal results are obtained when the patched model is compared with DPO on the generation of the patched model. This is intuitive, as it involves the actual generation-time activations of the patched model.
D.2 Which Token Position Should be Compared?
Previous studies typically utilized neuron activations at prompt tokens. We employed these activations to identify safety neurons for comparison. The results in Figure 16 and Table 10 indicate that safety neurons identified using generation-time activations yield more stable performance. However, Gemma-7b exhibits an unexpected behavior possibly due to the significantly different model architecture. We leave the investigation for the impact of model architectures on neuron-finding in future research.
Appendix E More Experimental Results
E.1 Correlation between GPT-4 Scores and Cost Scores
Evaluation with GPT-4 (Achiam et al., 2023) is also a widely accepted method (Liu et al., 2023; Dai et al., 2024). We leverage gpt-4-turbo-2024-04-09 to assign scores for the same generations from LLMs. The correlation between GPT-4 scores and cost scores is shown in Figure 9. We find there is a strong negative correlation between these two scores (-0.77), which indicates cost score is an appropriate metric for safety evaluation. The prompt and response of GPT-4 are demonstrated in Table 7.
Figure 9: The cost scores ($\downarrow$) and GPT-4 scores ($\uparrow$) of Llama2-7b SFT evaluate on Beavetrails. A strong negative correlation (-0.77) validates the effectiveness of cost scores as a faithful metric.
@@ E.2 Evaluation of Aligned Models
The average cost scores from the cost model are shown in Figure 10. Firstly, we noticed the models that have better performance in reports also perform better in safety alignment. Secondly, we find although SFT exhibit safety behaviors on average (due to the safety responses in ShareGPT), they are still vulnerable compared to DPO models. Thirdly, even if (IA)3 use only 0.005% parameters compared to full fine-tuning, it achieves relatively strong results in safety alignment (as a comparison, Llama2-7b-chat scores $-13.97$).
Figure 10: Cost scores of different models (lower is safer) evaluated on Beavertails.
E.3 More Alignment Tax Results

### v2

@@ 7 Related work
Neuron-Level Interpretability for Transformer. Identifying interpretable neurons has been a key goal in mechanistic interpretability for Transformers (Geva et al., 2021; Elhage et al., 2022; Gurnee et al., 2023, 2024). Geva et al. (2021) proposed viewing the feed-forward networks as key-value memories, offering a novel interpretive framework. Dai et al. (2022) identified knowledge neurons, whose activations correlate with facts, while Wang et al. (2022b) found skill neurons predictive of task labels. Gurnee et al. (2023) localized neurons relevant to specific features using sparse probing. However, these approaches are task-specific and not directly applicable to safety alignment. Gurnee et al. (2024) circumvented the need for ground-truth labels by using unsupervised methods to find universal neurons, leading to interpretable neuron families. Recent work by Lee et al. (2024) interpreted DPO in GPT-2 and identified toxic neurons affecting model behavior, while Yang et al. (2024) showed that DPO’s role extends beyond mitigating toxicity. Stolfo et al. (2024) identified confidence regulation neurons, revealing how induction heads use entropy neurons to adjust confidence. For safety-related neurons, predicting mechanistic patterns remains challenging.
Understanding Safety Mechanism of LLMs. Existing interpretability research on LLM safety can be broadly categorized into two perspectives: Representation Engineering (RepE, Zou et al., 2023) and Mechanistic Interpretability (MI, Elhage et al., 2021). RepE perspective takes a top-down approach, analyzing the residual stream to identify features (Zou et al., 2023; Zheng et al., 2024), which are then linked to neurons (Lee et al., 2024) or attention heads (Arditi et al., 2024). However, it is difficult to provide fine-grained mechanistic understanding and is usually used for steering model behavior rather than as an interpretability tool (Højer et al., 2025). MI perspective, on the other hand, takes a bottom-up approach, focusing on how basic model components contribute to safety. Wei et al. (2024) first introduced safety neurons as individual parameters. Since features in transformers are usually represented as vectors, it is difficult to interpret how different parameters in a single vector play different mechanistic roles. Li et al. (2025) adopted a safety layer perspective, which is too coarse-grained compared to neurons and attention heads for providing a mechanistic understanding. Zhou et al. (2025) identified key attention heads whose ablation increases attack success rates (ASR). Since MLP neurons constitute approximately two-thirds of the model’s parameters and serve as the foundational units for functionality, it is important to study safety mechanisms from the perspective of MLP neurons and further investigate their interaction with attention heads. A prior work (Zhao et al., 2025) identified safety neurons based on their contribution to the residual stream of safety-related data. This ablation-based method may be influenced by the existence of the "Hydra effect" (McGrath et al., 2023), which suggests that ablating certain components in LLMs may trigger compensatory behaviors in other components. Therefore, we adopted an opposite approach, enhancing rather than ablating, to locate neurons that have a causal effect on safety.
@@ 8 Conclusion
Acknowledgments
This work is supported by Beijing Natural Science Foundation (L243006) and National Natural Science Foundation of China (62476150).
Appendices
@@ A.3 Evaluation Benchmarks
TruthfulQA (Lin et al., 2022)
is a benchmark to measure whether a language model is truthful in generating answers to questions.
GSM8K (Grade School Math 8K, Cobbe et al., 2021)
is a dataset of 8.5K high-quality linguistically diverse grade school math word problems.
MMLU (Massive Multitask Language Understanding, Hendrycks et al., 2021)
is a massive multitask test consisting of multiple-choice questions from various branches of knowledge.
BBH (BIG Bench Hard, Suzgun et al., 2023)
is a subset of BIG Bench dataset and consists of 23 tasks that are particularly hard for the current generation of language models.
@@
RedTeam - $38,961$
HarmBench - $400$
JailbreakLLMs - $390$
LIMA - $1,030$
@@
MMLU - $14,042$
GSM8K $7473$ $1319$
TruthfulQA - $817$
BBH - $6511$
@@ B.1 Safety Alignment
Table 6: Hyperparameter of LLMs studied.
Model $d_{\mathrm{vocab}}$ $d_{\mathrm{model}}$ $d_{\mathrm{mlp}}$ $n_{\mathrm{layers}}$ $n_{\mathrm{heads}}$ #Neurons Activation
@@
Qwen2.5-3B $151,936$ $2,048$ $11,008$ $36$ $16$ $396,288$ SiLU
DPO Training Details
Details of (IA)3
Short for Invertible Adapters with Activation Alignment (Liu et al., 2022), (IA)3 is a fine-tuning method designed for large neural networks that achieves efficiency by focusing on a small number of trainable parameters while preserving the original model’s capacity. In our framework, we only apply (IA)3 to MLP as follows:
$\small\mathtt{MLP}(x)=\mathrm{W_{down}^{\top}}(\sigma(\mathrm{W_{gate}}~x)\odot\mathrm{W_{up}}~x\odot l_{\mathrm{ff}})$ (5)
where $l_{\mathrm{ff}}\in\mathbb{R}^{d_{m}}$ is the trainable parameters.
B.2 Evaluation Details
For general capabilities, we evaluate perplexity on the full test set of Wikitext-2 with a maximum length of 4096 and follow the evaluation settings outlined in Wang et al. (2024) for other benchmarks. Specifically, for MMLU, we use the entire test set and employ 0-shot prompting without Chain of Thought (CoT), selecting the option with the highest probability as the predicted choice, rather than using the model to generate the response directly. This approach differs from the method used in the official technical reports of these models, leading to some discrepancies in the results. For BBH, we sampled 40 samples from each task for testing and used a 3-shot CoT. For GSM8K, we sampled 200 samples using 8-shot CoT. For TruthfulQA, we utilize the official evaluation script, testing on the entire test set with the MC1 metric as proposed in Lin et al. (2022). The sampling strategy is the same as described before.
We run all the above experiments on NVIDIA A100-SXM4-80GB GPU, and it takes about 1,000 GPU hours.
B.3 Finding Safety Neurons
We build our code on TransformerLens (Nanda and Bloom, 2022) to cache neuron activations and perform dynamic activation patching. For each prompt dataset, we use 200 randomly sampled prompts (no overlap with evaluation data). Again, we use greedy search for generation and set the max new tokens to 256, resulting in around 40,000 activations for each neuron. We describe our dynamic activation patching method in Algorithm 1.
For the two ablation-based baselines, we referred to their official code and implemented our own version. For Wei et al. (2024), we used the Wanda score (Sun et al., 2024) as a parameter importance indicator to locate the top 5% of parameters in the down_proj matrix of each MLP layer in the DPO model, using the HH-Harmless dataset. For Zhao et al. (2025), we used the circuit breakers (Zou et al., 2025) training set, as described in their paper, to locate safety neurons, again selecting the top 5% as a comparison.
Algorithm 1 Dynamic Activation Patching
Input:
$w$ the prompt text
$\mathcal{N}$ a set contains (layer, neurons) pairs
$\mathcal{M}_{1}$ the model being patched
$\mathcal{M}_{2}$ the model used for patching
Output:
$w^{\prime}$ the completed text
Initialization:
   $w^{\prime}\leftarrow w$
   $l\leftarrow\mathcal{M}_{1}$.num_layers
repeat
[…]
  $x\leftarrow\mathcal{M}_{1}$.Embed($w^{\prime}$)
  for $i\leftarrow 1$ to $l$ do
   $x\leftarrow x+\mathcal{M}_{1}$.Attn[i]($x$)
   if $i$ in $\mathcal{N}$ then
[…]
   else
    $x\leftarrow x+\mathcal{M}_{1}$.MLP[i]$(x)$
   end if
  end for
  $p\leftarrow\mathcal{M}_{1}$.lm_head(x)[1](x)[-1].softmax()
  $token\leftarrowSample$$(p)$ {Get next token prediction}
  $w^{\prime}\leftarrowConcat$$(w^{\prime},token)$
until $\texttt{StopCriterion}(w^{\prime})$ is truetrue
B.4 Harmful Content Prediction
Appendix C More Properties of Safety Neurons
C.1 Layer Distribution
The layer distribution of the top $20,000$ safety neurons is shown in Figure 5(b). Llama2-7b and Mistral-7b have similar patterns: safety neurons are distributed across many layers, predominantly appearing in the deep layers, with a gradual shift towards the middle layers as change scores decrease. Conversely, Gemma-7b presents a starkly different distribution, with safety neurons primarily found in the initial and final layers. Notably, the most significant neurons in Gemma-7b are located in shallower layers, progressively transitioning to deeper layers with a more uniform distribution as change scores decrease. This phenomenon is likely due to significant architectural differences between Gemma-7b and the other two models (Table 6).
Figure 5: (a) The distribution of change scores of (20,000) safety neurons (truncated for better visualization). (b) The layer distribution of (20,000) safety neurons, grouped by every 5,000 neurons. The layer depth is the normalized layer number.
C.2 Change Score Distribution
We visualize the change scores distribution of top $20,000$ safety neurons in Figure 5(a). We first notice that only a small fraction of neurons changed much after safety alignment (for Llama2-7b only 876 out of 341248 neurons with a change score larger than $0.1$). More interestingly, these three different models have similar patterns and thresholds at around $0.035$ for safety neurons. Furthermore, we find that models performing better in safety alignment exhibit longer tails [^2] The skewness of Llama2-7b, Mistral-7b-v0.1 and Gemma-7b are $6.99$, $7.20$ and $19.89$ respectively., indicating that improved model performance may result from more neurons experiencing significant activation changes. We leave the further investigation of this phenomenon for future work.
C.3 Effectiveness of Change Score
We futher conducted experiments to validate whether the change score serves as an appropriate indicator of a neuron’s causal effect on generation. Specifically, we utilized consecutive sets of $5\%$ of neurons, starting from various ranks. As shown in Figure 6, we observed that as the change scores of the neurons decreased, the effectiveness of dynamic activation patching rapidly diminished. This finding indicates that only neurons with high change scores exert a significant causal effect on the model’s output. Consequently, we selected the top $5\%$ of neurons with the highest change scores as the safety neurons for further investigation in subsequent experiments.
Figure 6: Causal effects of different consecutive $5\%$ neurons in Base and SFT. The horizontal axis represents the rank of the highest-ranked neuron among these $5\%$ neurons (i.e., 0 refers to the safety neurons).
C.4 Specificity on Different Datasets
We simply use safety neurons found on HH-Harmless in previous experiments. Now we take a closer look at the prompt dataset selection. We use datasets from 3 different preference learning tasks: (1) Safety, including Beavertails (Ji et al., 2024), HH-Harmless (Bai et al., 2022a), and JailBreakLLMs (Shen et al., 2023); (2) Helpfulness, including HH-Harmless (Bai et al., 2022a) and LIMA (Zhou et al., 2024); (3) Reasoning, including the Reasoning subset from RewardBench (Lambert et al., 2024). We repeat the experiments from § 4.1 using safety neurons found on these prompts, as shown in Figure 7. The results indicate that safety neuron activations are specific to certain inputs, i.e., safety neurons found on similar types of prompts exhibit similar causal effects and are most effective on safety-related prompts.
Appendix D Other Design Choices for Neuron-Finding
D.1 Which Model Should be Compared?
We explore the impact of comparing different models and different generations. We replicate the experiments from § 4.1 with different design choices, and the results are depicted in Figure 8. These results indicate that there is no fundamental difference among the models chosen for comparison within our framework. However, the neurons identified by comparing SFT and DPO perform slightly better, which may be attributed to the minimal functional discrepancies between them, providing a clearer signal for identifying safety neurons.
D.2 Which Token Position Should be Compared?
Previous studies typically investigated neuron activations at prompt tokens (Zou et al., 2023). We employed these activations to identify safety neurons for comparison. The results in Figure 9 indicate that safety neurons identified using inference-time activations yield more stable performance. However, Gemma-7b exhibits an unexpected behavior possibly due to the significantly different model architecture. We leave the investigation for the impact of model architectures on neuron-finding in future research.
Appendix E More Experimental Results
E.1 Correlation between GPT-4 Scores and Cost Scores
Evaluation with GPT-4 (Achiam et al., 2023) is also a widely accepted metric (Liu et al., 2023; Dai et al., 2024). We leverage gpt-4-turbo-2024-04-09 to assign scores for the same generations from LLMs. The correlation between GPT-4 scores and cost scores is shown in Figure 10(a). We find there is a strong negative correlation between these two scores (-0.77), which indicates cost score is an appropriate metric for safety evaluation. The prompt and response of GPT-4 are demonstrated in Table 7.
Figure 10: (a) The cost scores ($\downarrow$) and GPT-4 scores ($\uparrow$) of Llama2 SFT evaluate on Beavetrails. A strong negative correlation (-0.77) validates the effectiveness of cost scores as a faithful metric. (b) The average accuracy of Llama2 classifier using different numbers of neuron activations.
@@ E.2 Evaluation of Aligned Models
The average cost scores of our SFT and DPO models on Beavertails can be found in Table 8. Firstly, we noticed the models that have better performance in reports also perform better in safety alignment. Secondly, we find although SFT exhibit safety behaviors on average (due to the safety responses in ShareGPT), they are still vulnerable compared to DPO models. Thirdly, even if (IA)3 use only 0.005% parameters compared to full fine-tuning, it achieves relatively strong results in safety alignment (as a comparison, Llama2-7b-chat scores 13.97-13.97).
E.3 More Safety Neuron Results
In Table 8, we present the complete results of the cost scores and general benchmark performance.
Table 8: Cost scores on red-teaming benchmarks and general capabilities on various benchmarks. Abbr. BT = Beavertails, RT = RedTeam, HB = HarmBench, JL = JailBreakLLMs, GSM = GSM8K, TQA = TruthfulQA. / / § denotes patching with neurons identified by Param (Wei et al., 2024) / Neuron (Zhao et al., 2025) / Ours. Bold denotes the best performance within the same model class.
Model BT ($\downarrow$) RT ($\downarrow$) HB ($\downarrow$) JL ($\downarrow$) PPL ($\downarrow$) GSM ($\uparrow$) BBH ($\uparrow$) MMLU ($\uparrow$) TQA ($\uparrow$)
Llama2 Base $2.2$ $5.7$ $8.0$ $1.1$ $\mathbf{5.1}$ $0.150$ $0.139$ $\mathbf{0.398}$ $0.252$
Base $1.6$ $5.9$ $8.1$ $1.2$ $5.1$ $\mathbf{0.165}$ $\mathbf{0.141}$ $0.395$ $0.248$
Base $2.6$ $5.6$ $8.9$ $0.5$ $5.2$ $0.140$ $0.131$ $0.387$ $\mathbf{0.259}$
Base§ $\mathbf{-5.7}$ $\mathbf{-5.7}$ $\mathbf{-3.9}$ $\mathbf{-7.9}$ $5.6$ $0.100$ $0.131$ $0.392$ $0.257$
SFT 2.4-2.4 2.9-2.9 $5.0$ $4.0$ $5.4$ $0.095$ $0.110$ $0.398$ $0.263$
SFT 2.7-2.7 2.8-2.8 $4.3$ $4.1$ $5.4$ $0.100$ $0.109$ $0.391$ $0.261$
SFT 3.6-3.6 5.6-5.6 $3.3$ $2.9$ $5.4$ $0.080$ $0.110$ $0.389$ $0.267$
SFT§ $\mathbf{-11.9}$ $\mathbf{-11.9}$ $\mathbf{-7.2}$ $\mathbf{-6.6}$ $\mathbf{5.4}$ $\mathbf{0.105}$ $\mathbf{0.131}$ $\mathbf{0.399}$ $\mathbf{0.277}$
DPO 11.8-11.8 11.8-11.8 11.0-11.0 10.5-10.5 $5.5$ $0.095$ $0.094$ $0.374$ $0.280$
Mistral Base 1.6-1.6 4.8-4.8 1.1-1.1 $3.2$ $\mathbf{4.9}$ $\mathbf{0.285}$ $\mathbf{0.169}$ $\mathbf{0.578}$ $0.284$
Base 1.1-1.1 5.1-5.1 1.3-1.3 $2.8$ $4.9$ $0.265$ $0.165$ $0.578$ $0.283$
Base 1.6-1.6 5.8-5.8 0.9-0.9 1.1-1.1 $5.1$ $0.225$ $0.169$ $0.543$ $0.272$
Base§ $\mathbf{-10.0}$ $\mathbf{-10.2}$ $\mathbf{-7.8}$ $\mathbf{-8.5}$ $5.1$ $0.125$ $0.163$ $0.573$ $\mathbf{0.296}$
SFT 7.6-7.6 7.3-7.3 $3.7$ $0.2$ $\mathbf{5.2}$ $0.215$ $0.168$ $\mathbf{0.583}$ $0.275$
SFT 7.3-7.3 7.7-7.7 $4.1$ $0.7$ $5.2$ $0.230$ $\mathbf{0.177}$ $0.583$ $0.278$
SFT 8.2-8.2 9.1-9.1 $1.6$ 0.8-0.8 $5.2$ $0.240$ $0.169$ $0.570$ $0.275$
SFT§ $\mathbf{-12.9}$ $\mathbf{-12.2}$ $\mathbf{-3.6}$ $\mathbf{-6.1}$ $5.3$ $\mathbf{0.265}$ $0.170$ $0.579$ $\mathbf{0.282}$
DPO 13.5-13.5 13.4-13.4 6.1-6.1 8.2-8.2 $5.3$ $0.140$ $0.163$ $0.576$ $0.288$
Gemma Base $1.1$ $0.4$ $7.8$ $1.1$ $\mathbf{6.6}$ $0.080$ $0.223$ $0.599$ $0.311$
Base $0.8$ $1.8$ $6.7$ $0.6$ $6.6$ $0.080$ $\mathbf{0.224}$ $\mathbf{0.600}$ $\mathbf{0.312}$
Base $8.5$ $8.0$ $8.5$ $7.5$ $8.1$ $0.090$ $0.188$ $0.531$ $0.297$
Base§ $\mathbf{-10.3}$ $\mathbf{-9.5}$ $\mathbf{-4.8}$ $\mathbf{-7.1}$ $7.0$ $\mathbf{0.100}$ $0.208$ $0.578$ $0.301$
SFT 8.2-8.2 9.8-9.8 $1.0$ 1.6-1.6 $\mathbf{7.5}$ $0.345$ $\mathbf{0.217}$ $0.571$ $0.321$
SFT 8.7-8.7 10.4-10.4 $1.3$ 0.9-0.9 $7.5$ $\mathbf{0.350}$ $0.214$ $0.574$ $\mathbf{0.322}$
SFT 7.7-7.7 9.0-9.0 $4.3$ 0.3-0.3 $7.7$ $0.320$ $0.210$ $\mathbf{0.588}$ $0.315$
SFT§ $\mathbf{-13.4}$ $\mathbf{-13.4}$ $\mathbf{-9.2}$ $\mathbf{-9.6}$ $7.6$ $0.300$ $0.213$ $0.565$ $0.312$
DPO 13.6-13.6 14.1-14.1 11.9-11.9 10.6-10.6 $7.9$ $0.200$ $0.196$ $0.549$ $0.324$
Qwen2.5 Base $0.9$ 1.3-1.3 $3.7$ $6.9$ $\mathbf{7.6}$ $\mathbf{0.355}$ $\mathbf{0.185}$ $\mathbf{0.646}$ $0.334$
Base $1.1$ 1.3-1.3 $3.8$ $6.0$ $7.6$ $0.375$ $0.179$ $0.646$ $0.332$
Base 3.5-3.5 2.1-2.1 $3.3$ $4.5$ $7.8$ $0.335$ $0.169$ $0.638$ $0.334$
Base§ $\mathbf{-12.0}$ $\mathbf{-12.4}$ $\mathbf{-8.1}$ $\mathbf{-8.8}$ $7.9$ $0.535$ $0.198$ $0.644$ $\mathbf{0.343}$
SFT 10.9-10.9 9.9-9.9 2.8-2.8 2.3-2.3 $\mathbf{7.9}$ $0.590$ $0.171$ $\mathbf{0.635}$ $0.344$
SFT 10.9-10.9 9.9-9.9 3.1-3.1 2.3-2.3 $7.8$ $0.595$ $\mathbf{0.167}$ $0.635$ $0.341$
SFT 10.8-10.8 9.7-9.7 3.1-3.1 2.1-2.1 $7.9$ $0.585$ $0.168$ $0.636$ $0.346$
SFT§ $\mathbf{-13.3}$ $\mathbf{-13.3}$ $\mathbf{-8.2}$ $\mathbf{-8.7}$ $7.9$ $\mathbf{0.600}$ $0.175$ $0.636$ $\mathbf{0.349}$
DPO 13.8-13.8 14.1-14.1 10.7-10.7 10.0-10.0 $8.0$ $0.560$ $0.167$ $0.632$ $0.354$
In Table 9, we present the complete results of the top safety neurons’ value vectors projected into the vocabulary space.
Table 9: Top safety neuron value vectors from Llama2-7b projected onto the vocabulary space. MLP.vllnn    denotes the down projection vector of the $n$-th neuron in layer $l$. We omitted some tokens for better visualization.
Vector Top Tokens
MLP.v3010106 ouc, iter, trat, ussen, tid, imos
MLP.v298343 </s>, Genomsnittlig, ←, text, <s>
MLP.v285293 Sug, Commons, sugar, mouth, flesh
[…]
MLP.v267581 wa, ales, sin, MainActivity, oblig
MLP.v299647 Food, Guard, Farm, Ali, Sex, Break
[…]
MLP.v284127 **, >>.***, °, ’’’, ----, /, !!, ]
MLP.v307219 Ż, Gemeinsame, HT, Gor, category
E.4 More Alignment Tax Results